of 30 students in a class, 25 speak french and 20 speak english. each student speaks atleast 1 of these languages.
we meet a student from the class at random.
what is the probability that this student speaks:
1-both languages.
2-french only.
3-one language only.

Answer by ewatrrr(24785)

 
Hi,
In reply to Your TY: 45 people speak one or both of the languages...
15 must speak both, because there are only 30 people: 45-30 = 15
30 students in a class: 25 speak french and 20 speak english.
each student speaks atleast 1 of these languages
P(both languages) = 15/30 = 1/2
P(french only) = 10/30 = 1/3
P(one language only) = 15/30 = 1/2